# Generating CAD Code with Vision-Language Models for 3D Designs

### v1

@@ Abstract
Generative AI has transformed the fields of Design and Manufacturing by providing efficient and automated methods for generating and modifying 3D objects. One approach involves using Large Language Models (LLMs) to generate Computer-Aided Design (CAD) scripting code, which can then be executed to render a 3D object; however, the resulting 3D object may not meet the specified requirements. Testing the correctness of CAD generated code is challenging due to the complexity and structure of 3D objects (e.g., shapes, surfaces, and dimensions) that are not feasible in code. In this paper, we introduce CADCodeVerify, a novel approach to iteratively verify and improve 3D objects generated from CAD code. Our approach works by producing ameliorative feedback by prompting a Vision-Language Model (VLM) to generate and answer a set of validation questions to verify the generated object and prompt the VLM to correct deviations. To evaluate CADCodeVerify, we introduce, CADPrompt, the first benchmark for CAD code generation, consisting of $200$ natural language prompts paired with expert-annotated scripting code for 3D objects to benchmark progress. Our findings show that CADCodeVerify improves VLM performance by providing visual feedback, enhancing the structure of the 3D objects, and increasing the success rate of the compiled program. When applied to GPT-4, CADCodeVerify achieved a 7.30% reduction in Point Cloud distance and a 5.0% improvement in success rate compared to prior work.
@@ 1 Introduction
Generative AI, such as Large Language Models (LLMs) and Vision-Language Models (VLMs), offers a unique opportunity to enhance productivity, reduce costs, and increase efficiency within Design and Manufacturing sectors  (Kumar et al., 2023). These industries are critical contributors to the global economy and responsible for creating products and infrastructure. Recent work has shown that Generative AI can provide solutions towards generating, evaluating, and correcting designs of 3D objects  (Nelson et al., 2023; Kodnongbua et al., 2023; Makatura et al., 2023). However, while these solutions improve efficiency for designers and engineers, they lack efficient feedback mechanisms or refinement loops to automatically correct any inaccuracies in the initially generated 3D object.
@@
Our research explores the potential of using VLMs to generate and refine Computer-Aided Designs (CAD), the leading method in industrial design for creating, modifying, analyzing, and optimizing 3D objects (Sarcar et al., 2008). Utilizing CAD software such as FreeCAD (Riegel et al., 2016) and AutoCAD (Yarwood, 2013) to design products typically requires significant training and domain expertise. CAD software typically includes underlying scripting languages which allow a user to build parametric 3D objects using code, rather than through the user interface. Leveraging the code generating faculties of Generative AI allows us to circumvent the challenges of utilizing CAD software by directly generating the underlying scripting code representing a 3D object.
@@
We define the process of generating and refining 3D objects using LLMs or VLMs, via CAD scripting code, as CAD code generation. Designs produced by off-the-shelf VLMs or LLMs frequently deviate functionally or structurally from stakeholder specifications, because Generative AI models tend to hallucinate on complex, out-of-distribution tasks. Developing refinement methods to correct inaccuracies in a generated design is a critical next-step in this research. Current state-of-the-art refinement methods are contingent on human-in-the-loop domain expertise  (Makatura et al., 2023; Nelson et al., 2023). Obtaining a domain expert is not always feasible and can be time consuming. Automated refinement mechanisms are necessary to improve on the initially imperfect designs and close the feedback loop. Augmenting CAD code generation methods with fully-automated feedback mechanisms will empower non-expert end users to generate their own designs, and reduce the onerous, iterative design process.
@@
In this work, we introduce CADCodeVerify, a fully automated method for refining CAD code generation. CADCodeVerify eliminates the need for human involvement by generating and answering validation questions based on user requirements and providing feedback to refine 3D object code. This feedback is employed to iteratively refine the design. To evaluate CADCodeVerify, we introduce a novel benchmark, CADPrompt, designed specifically for assessing the performance of Generative AI (e.g., LLMs and VLMs) in CAD code generation and refinement. CADPrompt contains $200$ 3D objects annotated with natural language prompts and expert-written Python code. This multimodal benchmark enables researchers to evaluate aspects of CAD code generation such as object quality and the syntactical correctness of the generated code. Our experimental results demonstrate that CADCodeVerify not only enhances the quality of generated 3D objects but also improves the performance of LLMs and VLMs by increasing the success rate of compiled programs.
Our work presents three key contributions:
@@
    With our benchmark, CADPrompt, we provide the first quantitative evaluation of CAD code generation capabilities across widely used VLMs, including GPT-4, Gemini 1.5 Pro, and CodeLlama.
@@
    We demonstrate that CADCodeVerify sets a new state-of-the-art in refining 3D designs via CAD scripting, achieving a $7.30\%$ reduction in Point Cloud distance and a $5.5\%$ increase in successful object generation when applied to GPT-4, the leading VLM for CAD code generation.
Figure 1: Our approach enables VLMs to automatically generate and refine 3D objects via a CAD scripting code (i.e., CADQuery) in three steps; (1) Code Generation, where the VLM generates CAD scripting code from a language prompt, (2) Code Execution, where the code generated by the model is rendered as a 3D object through a compiler, and (3) Code Refinement, wherein the language model engages in a self-initiated question-answering process to validate the generated object, with respect to the initial prompt, to generate actionable feedback to refine the code.
@@ 2 Related Work
We review two integral categories of prior work: LLMs for code generation and recent advances in automated design and manufacturing.
@@ 2.1 LLMs for Code Generation
LLMs have demonstrated impressive results in a variety of code generation applications. The most popular application is the text-to-code task, wherein users prompt LLMs with a description of the code, which facilitates the LLM to generate the code (Rozière et al., 2023; Svyatkovskiy et al., 2020; Brown et al., 2020; Poesia et al., 2022). Another application is the text-to-SQL task, in which a database and question are provided to the LLMs, and then the LLMs generate the SQL query (Rajkumar et al., 2022). We build upon these contributions to explore how LLMs can assist non-experts and designers to generate 3D objects via CAD code.
Prior research has demonstrated that LLMs can notably improve their generated code through feedback pertaining to the quality of the generated code  (Chen et al., 2023a). One approach involves LLMs providing their own feedback, wherein the LLMs generate explanatory rationales for the initially generated code. These rationales are then used as feedback, along with the initial code, to refine the output (Chen et al., 2023b). Another approach involves receiving feedback from external tools, such as static code analysis (Alrashedy et al., 2024), or a Python Interpreter (Madaan et al., 2023). Contrary to other works, self-correction  (Welleck et al., 2022), involves training two models: a generator and a corrector. The generator model produces the initial output, which is then refined by the corrector model. In CAD code refinement, one line of work introduced a basic feedback mechanism  (Yuan et al., 2024), using VLMs to refine 3D object code based on a script and an image. However, this method still struggles with feedback efficiency and refinement detail. In this work, we propose a novel approach to generate visual feedback that enhances the shapes, surfaces, and dimensions of 3D objects through code refinement.
@@ 2.2 Language for Automated Design and Manufacturing
Learning a shared embedding space between language and 3D objects is essential to integrate language modeling into design and manufacturing. Recent work has developed foundational multimodal transformers in the image-text space (Radford et al., 2021; Li et al., 2023a; Liu et al., 2023), which can be leveraged to perform a variety of tasks. These approaches have inspired similar techniques to integrating 3D objects into this a shared embedding space (Zeng et al., 2023; Xue et al., 2023; Yu et al., 2022; Guo et al., 2023), thereby facilitating a range of multimodal tasks such as dialog, classification and evaluation with 3D data (Hong et al., 2023; Xu et al., 2023). While these approaches are capable of learning a shared embedding space between language and 3D objects, they are incapable of generating 3D objects.
Generative modeling of 3D data has historically relied on probabilistic models for 3D space, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Variational AutoEncoders (VAEs) (Kingma & Welling, 2013), and Diffusion Models (Ho et al., 2020). While GANs and VAEs (Wu et al., 2015; Achlioptas et al., 2018; Gadelha et al., 2018; Yang et al., 2019; Yan et al., 2024) have historically been popular approaches, diffusion models have recently emerged as the state-of-the-art for probabilistic 3D modeling (Zeng et al., 2022; Koo et al., 2023). Diffusion models facilitate controllable object generation through language-guided shape generation and completion albeit their applicability in producing manufacturable designs remains limited  (Li et al., 2023b; Kodnongbua et al., 2023). These models typically output point clouds or voxels, neither of which are readily adaptable to manufacturing, due to their non-parametric nature. Thus, methodologies producing parametric outputs suitable for manufacturing are needed.
Recent studies have revealed the capability of VLMs to produce designs via parametric CAD code (Makatura et al., 2023; Nelson et al., 2023); however, these works find that GPT-4 requires a significant amount of human feedback to produce code that accurately reflects the initial specifications. In this paper, we address this limitation by developing CADCodeVerify to alleviate the need for a human expert by allowing a VLM to automatically generate and adapt a 3D object.
@@ 3 CAD Code Generation
Our goal is to enable VLMs to accurately generate 3D objects from non-expert language specifications. As shown in Figure 1, we take a three-step approach: (1) code generation (§3.1), (2) code execution (§3.2), and (3) code refinement via CADCodeVerify (§3.3).
@@ 3.1 Code Generation
We prompt the VLM to generate initial CAD scripting code, $y_{0}$, based on a natural language description of the 3D object, $x$, along with task specifications. For the few-shot experiment, we include a set of few-shot demonstrations, $E_{g}$, in the prompt (see Appendix B.4). We employ CADQuery, a Python-based parametric CAD language equipped with a compiler that processes code, to output parametric 3D objects, i.e., render.
Thus, CAD code generation can be broadly formulated as per Equation 1.
@@ 3.2 Code Execution
The generated CAD code, $y_{0}$, is next executed by the CADQuery compiler, $\psi$, to produce the initial CAD design, $d_{0}$, in the Standard Triangle Language (STL) format, [^1] Standard Triangle Language is also referred to as “Standard Tessellation Language”. The STL (STereoLithography) file format is an openly documented format for describing the surface of an object as a triangular mesh, that is, as a representation of a 3-dimensional surface in triangular facets. where 00 indicates the initial version of both the generated code and CAD design. We denote this process as $d_{0}\sim\psi(y_{0})$. Occasionally, the model produces code which does not compile, typically caused by syntactical issues in the Python code. We apply insights from prior work on code repair (Chen et al., 2023b), to CAD code generation. To resolve syntactical issues in the code, we leverage the CADQuery compiler error as feedback, $F_{err}$, for the VLM (see Appendix Figure 8, for the prompt employed in code repair). This process is repeated until a 3D-object, $d_{k}$, is successfully rendered by $\psi(y_{k})$, or the maximum number of iterations $N$ is reached as per Equation 2.
@@ 3.3 Code Refinement via CADCodeVerify
To correct any discrepencies in the generated design, we construct feedback, $F_{ref}$, to further refine $y_{k}$ as described by Equation 3, where $M$ is the number of refinement iterations, [^2] In all our experiments, we set the number of refinements to 2, as we did not observe any improvement after the second refinement, which is also consistent with prior work on refinement for code generation (Madaan et al., 2023; Chen et al., 2023a). We optionally include a set of four reference images of the generated object, taken from different angles (0, 90, 180, and 270 degrees), $I_{ref}=\{I_{l}^{0},I_{l}^{90},I_{l}^{180},I_{l}^{270}\}$.
@@
The key intuition behind our approach, CADCodeVerify, is to develop a feedback mechanism that leverages a vision-language model to produce visual feedback, which helps improve the generated 3D object without needing human expertise as in prior work (Makatura et al., 2023) or any external tools (e.g., geometric solvers — see §5.1). CADCodeVerify consists of two sub-steps:
@@
CADCodeVerify generates a set of binary Yes/No verification questions, $\mathrm{Q}=\{q_{1},q_{2},\dots,q_{n}\}$, based on the language description, $x$, and a set of few-shot example questions, $E_{q}$, i.e., $\mathrm{Q}\sim P_{LM}(\mathrm{Q}|x,E_{q})$. CADCodeVerify generates between two to five questions per example (see prompt in Figure 7 in the Appendix). To generate answers, CADCodeVerify uses images of the generated 3D object ${I_{ref}}$, the generated questions $Q$, and the language description $x$, where CADCodeVerify answers each of the validation questions as described in Equation 4.
@@
CADCodeVerify generates the answers, $\mathrm{A}=(a_{1},a_{2},\dots,a_{n})$, using the Chain-of-Thought format (Wei et al., 2022), where each answer is accompanied by corresponding reasoning (see prompt example in Figure 9 in the Appendix). Additionally, CADCodeVerify is instructed to respond with "Unclear" when it determines that sufficient information is lacking to accurately answer the question.
@@
The question-answer sets are next used to generate ameliorative feedback for refining the 3D object as per Equation 5.
@@
The feedback, $F_{ref}$, is then utilized for the code refinement, as described Equation (3). During the feedback generation step, we exclude any questions for which the answer is “Yes” to enable the model to focus on resolving any outstanding issues. If the answer to all questions is “Yes,” then we assume that no further refinement is needed. The full feedback generation prompt can be found in Figure 10 in the Appendix. Two examples of the complete interactions with GPT-4 for one step of CADCodeVerify have been included in the Appendix (see Figures 5 and 6).
Figure 2: Examples of 3D objects generated by GPT4 after utilizing each feedback method for code refinement. Column 1 shows the ground truth object. The second column shows the language prompt passed into the vision-language model. Columns three, four, and five show the generated object after the initial generation, first-refine, and second-refine phases. In each refine phase, we show the object generated by each feedback method used for code refinement. The X icon denotes situations wherein GPT4 generated code did not compile.
Table 1: Statistics of the CADPrompt dataset, including the vertex and face counts of ground truth 3D objects, the lengths of language descriptions, as well as Python code. We split the dataset into simple and complex categories (§4.3).
Structural Complexity of 3D Object Natural Language Descriptions Python Code
Category Vertices Faces Words Sentences Lines of Code Tokens
@@
Simple 6 44 16.7 8 84 30.0 9 188 51.0 1 11 3.4 6 46 16.9 18 117 46.3
Complex 52 540 161.7 100 1092 323.2 9 152 49.5 1 10 3.3 6 40 16.3 18 104 43.6
All 6 540 88.8 8 1092 175.7 9 188 50.13 1 11 3.4 6 46 16.6 18 117 45.0

### v2

@@ Abstract
Generative AI has revolutionized the fields of Design and Manufacturing by providing efficient and automated methods for generating and modifying 3D objects. One approach involves using Large Language Models (LLMs) to generate Computer-Aided Design (CAD) scripting code, which can then be executed to render a 3D object; however, the resulting 3D object may not meet the specified requirements. Testing the correctness of CAD-generated code is challenging due to the structural intricacies of 3D objects that are not discernable in code. In this paper, we introduce CADCodeVerify, a novel approach to iteratively verify and improve the design output of 3D objects generated from CAD code. Our approach provides ameliorative feedback by prompting a Vision Language Model (VLM) to generate and answer a set of validation questions to verify the generated object and prompt the VLM to correct deviations. To evaluate CADCodeVerify, we introduce, CADPrompt, the first benchmark for CAD code generation, consisting of $200$ natural language prompts paired with expert-annotated scripting code for 3D objects to benchmark progress. Our findings show that CADCodeVerify improves VLM performance by providing visual feedback by enhancing the structure of the 3D objects and increasing the compile rate of the compiled program. When applied to GPT-4, CADCodeVerify achieved a 7.30% reduction in Point Cloud distance and a 5.5% improvement in compile rate compared to prior work. Code and data are available at https://github.com/Kamel773/CAD_Code_Generation
@@ 1 Introduction
Generative AI, such as Large Language Models (LLMs) offers a unique opportunity to enhance productivity, reduce costs, and increase efficiency within Design and Manufacturing sectors  (Kumar et al., 2023). These industries are critical contributors to the global economy and responsible for creating products and infrastructure. Recent research has demonstrated that Generative AI can support the generation, evaluation, and correction of 3D object designs (Nelson et al., 2023; Kodnongbua et al., 2023; Makatura et al., 2023). However, while these solutions improve efficiency for designers and engineers, they often lack effective feedback mechanisms or refinement loops to automatically address inaccuracies in the initially generated 3D object.
@@
Our research explores the potential of using VLMs to generate and refine Computer-Aided Designs (CAD), the CAD a leading approach in industrial design for creating, modifying, analyzing, and optimizing 3D objects (Sarcar et al., 2008). Designing products with CAD software such as FreeCAD (Riegel et al., 2016) and AutoCAD (Yarwood, 2013) generally requires substantial training and domain expertise. CAD software often includes scripting languages that allow users to build parametric 3D objects using code, rather than relying solely on the user interface. Leveraging the code-generating capabilities of Generative AI enables users to bypass the complexities of traditional CAD software by directly generating the underlying scripting code for representing 3D objects.
@@
We define the process of generating and refining 3D objects using LLMs or VLMs, through CAD scripting code, as CAD code generation. Designs produced by off-the-shelf VLMs or LLMs often deviate functionally or structurally from stakeholder specifications as these models often hallucinate on complex, out-of-distribution tasks. Developing refinement methods to correct inaccuracies in a generated design is a critical next step in this research. Current state-of-the-art refinement methods are contingent on human-in-the-loop expertise  (Makatura et al., 2023; Nelson et al., 2023), which can be time- and cost-prohibitive. Instead, we propose to develop automated feedback mechanisms for CAD code generation to reduce the barrier to entry for design and expedite lengthy design processes.
@@
In this work, we introduce CADCodeVerify, an automated method for refining CAD code generation. CADCodeVerify eliminates the need for human involvement by generating and answering validation questions based on user requirements, offering feedback to refine 3D object code. This feedback is used to iteratively refine the design. To evaluate the performance of CADCodeVerify, we introduce a novel benchmark, CADPrompt contains $200$ 3D objects annotated with natural language prompts and expert-written Python code. This multimodal benchmark enables researchers to evaluate aspects of CAD code generation, including object quality and the syntactical correctness of the generated code. Our experimental results demonstrate that CADCodeVerify not only enhances the quality of generated 3D objects but also improves the performance of LLMs and VLMs by increasing the compile rate of compiled programs. Our work presents three key contributions:
@@
    With our benchmark, CADPrompt, we provide the first quantitative evaluation of CAD code generation across GPT-4, Gemini 1.5 Pro, and CodeLlama.
@@
    We demonstrate that CADCodeVerify sets a new state-of-the-art 3D design via CAD scripting, achieving a $7.30\%$ reduction in Point Cloud distance and a $5.5\%$ increase in successful object generation using to GPT-4, the leading VLM for CAD code generation.
Figure 1: Our approach enables VLMs to automatically generate and refine 3D objects through a CAD scripting code (e.g., CADQuery) in three steps; (1) Code Generation, where the VLM generates CAD scripting code from a language prompt, (2) Code Execution, where the code generated by the model is rendered as a 3D object through a compiler, and (3) Code Refinement, wherein the language model engages in a self-initiated question-answering process to validate the generated object, with respect to the initial prompt, to generate actionable feedback to refine the code.
@@ 2.1 LLMs for Code Generation
LLMs have demonstrated impressive results in a variety of code generation applications. The most popular application is the text-to-code task, wherein users prompt LLMs with a code description, enabling the LLM to generate the code (Rozière et al., 2023; Svyatkovskiy et al., 2020; Brown et al., 2020; Poesia et al., 2022). Another application is the text-to-SQL task, where a database and a question are provided to the LLMs, which then generate the corresponding SQL query (Rajkumar et al., 2022). Building on these advances, we explore how LLMs can support non-experts and designers in generating 3D objects via CAD code.
Prior research has demonstrated that LLMs can significantly improve the quality their generated code by incorporating feedback  (Chen et al., 2023a). One such approach involves LLMs providing their own feedback by generating explanatory rationales for the initially generated code. These rationales are then used as feedback, along with the initial code, to refine the output (Chen et al., 2023b). Another approach relies on feedback from external tools, such as static code analysis (Alrashedy et al., 2024), or a Python Interpreter (Madaan et al., 2023).
In contrast, the self-correction method  (Welleck et al., 2022) trains two separate models: a generator, which creates the initial output, and a corrector, which refines the generated code. In CAD code refinement, one line of research introduces a basic feedback mechanism (Yuan et al., 2024), wherein VLMs refine 3D object code based both on a script and an image. However, this method still faces challenges in terms of feedback efficiency and refinement precision. In this work, we propose a novel approach to generating visual feedback that improves the shapes, surfaces, and dimensions of 3D objects through more effective code refinement.
@@ 2.2 Language for Automated Design and Manufacturing
Learning a shared embedding space between language and 3D objects is essential for integrating language into design and manufacturing. Recent advancements have introduced foundational multimodal transformers for image-text data  (Radford et al., 2021; Li et al., 2023a; Liu et al., 2023) which can be used to perform a variety of tasks. These techniques have inspired similar approaches for incorporating 3D objects into shared embedding spaces  (Zeng et al., 2023; Xue et al., 2023; Yu et al., 2022; Guo et al., 2023), enabling a wide range of multimodal tasks such as dialogue, classification, and evaluation involving 3D data (Hong et al., 2023; Xu et al., 2023). While successfully learning a shared-embedding space, these approaches are not capable of generating 3D objects.
Generative modeling of 3D data has traditionally relied on probabilistic models for 3D space. These include: Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Variational AutoEncoders (VAEs) (Kingma & Welling, 2013), and Diffusion Models (Ho et al., 2020). While GANs and VAEs (Wu et al., 2015; Achlioptas et al., 2018; Gadelha et al., 2018; Yang et al., 2019; Yan et al., 2024) have been popular approaches, diffusion models have recently emerged as the state-of-the-art for probabilistic 3D modeling (Zeng et al., 2022; Koo et al., 2023). Diffusion models enable controllable object generation through language-guided shape generation and completion. However, their application in producing manufacturable designs remains limited  (Li et al., 2023b; Kodnongbua et al., 2023). These models typically generate point clouds or voxels, which are non-parametric and not easily adapted for manufacturing. As such, there is a need for methods that produce parametric outputs suitable for manufacturing.
Recent research has shown that VLMs can generate designs using parametric CAD code (Makatura et al., 2023; Nelson et al., 2023). However, these approaches still require significant human feedback to produce code ensure the generated codes meets the user’s specifications. In this paper, we address this limitation by developing CADCodeVerify, which autonomously generates and refines 3D objects.
@@ 3 CAD Code Generation
As shown in Figure 1, we employ a three-step process: (1) code generation (§3.1), (2) code execution (§3.2), and (3) code refinement via CADCodeVerify for CAD code generation(§3.3).
@@ 3.1 Code Generation
We prompt the VLM to generate initial CAD script code, $y_{0}$, based on a natural language description of the 3D object, $x$, along with task specifications. For the few-shot experiment, we incorporate a set of few-shot demonstrations, $E_{g}$, in the prompt (see Appendix B.4). We utilize CADQuery, a Python-based parametric CAD language with a built-in compiler that interprets code and renders parametric 3D object. This CAD code generation can be broadly formulated as per Eq. 1.
@@ 3.2 Code Execution
The generated CAD code, $y_{0}$, is subsequently executed by the CADQuery compiler, $\psi$, to produce the initial CAD design, $d_{0}$, in the Standard Triangle Language (STL) format, [^1] Standard Triangle Language is also referred to as “Standard Tessellation Language”. The STL (STereoLithography) file format is an openly documented format for describing the surface of an object as a triangular mesh, that is, as a representation of a 3-dimensional surface in triangular facets. where 00 indicates the initial version of both the generated code and CAD design. We denote this process as $d_{0}=\psi(y_{0})$. Occasionally, the models generate code that fails to compile, usually due to syntax errors in the Python code. We leverage insights from previous work on code repair (Chen et al., 2023b) and adapt similar techniques for CAD code generation. To resolve syntactical issues in the code, we leverage the CADQuery compiler error as feedback, $F_{err}$, for the VLM (see Appendix Figure 12, for the prompt employed in code repair). This process is repeated until a 3D object, $d_{k}$, is successfully rendered by $\psi(y_{k})$, or the maximum number of iterations, $N$, is reached as per Eq. 2.
@@ 3.3 Code Refinement via CADCodeVerify
To address discrepancies in the generated design, we use a feedback loop to further refine the CAD code, $F_{ref}$, to further refine $y_{k}$ as described by Eq. 3, where $M$ is the number of refinement iterations. Optionally, we include a set of four reference images of the generated object, captured from different angles (0, 90, 180, and 270 degrees), $I_{ref}=\{I_{l}^{0},I_{l}^{90},I_{l}^{180},I_{l}^{270}\}$.
@@
The key innovation of CADCodeVerify is its feedback loop, which uses a VLM to generate visual feedback automatically, without human intervention or the need for external tools like geometric solvers  (Makatura et al., 2023) (see §5.1). CADCodeVerify is a two-step process consisting of question-answer generation and feedback generation.
@@
CADCodeVerify first generates a set of binary “Yes/No” verification questions, $\mathrm{Q}=\{q_{1},q_{2},\dots,q_{n}\}$, based on the language description, $x$, and a set of few-shot example questions, $E_{q}$, i.e., $\mathrm{Q}\sim P_{LM}(\mathrm{Q}|x,E_{q})$. CADCodeVerify generates between two to five questions per example (see Figure 11 in the Appendix). To answer these questions, CADCodeVerify uses the reference images of the generated 3D object, ${I_{ref}}$, the generated questions, $Q$, and the language description $x$. CADCodeVerify then answers each question as described in Eq. 4.
@@
CADCodeVerify generates the answers, $\mathrm{A}=(a_{1},a_{2},\dots,a_{n})$, using Chain-of-Thought (Wei et al., 2022), where each answer is accompanied by supporting reasoning (see Figure 13 in the Appendix). Furthermore, CADCodeVerify is directed to respond with "Unclear" if it determines that there is insufficient information to answer the question.
@@
The question-answer pairs are then used to generate ameliorative feedback to further refine the 3D object as per Eq. 5.
@@
feedback, $F_{ref}$, is then applied to refine the code, as described Eq. (3). During this step, we omit any questions for which the answer is “Yes” to allow the model to focus on addressing any unresolved issues. If all questions are answered “Yes” we assume no further refinement is necessary. The full feedback generation process is provided in Figure 14 in the Appendix. Additionally, two examples of full interactions with GPT-4 for a single iteration of CADCodeVerify are included in the Appendix ( Figures 9 and 10).
Figure 2: Examples of 3D objects generated by GPT4 after utilizing each feedback method for code refinement from our CADPrompt dataset.
Table 1: Statistics of the CADPrompt dataset, including the vertex and face counts of ground truth 3D objects, as well as the lengths of language descriptions and Python code (§4.3).
Structural Complexity of 3D Object Natural Language Descriptions Python Code Total
Category Vertices Faces Words Sentences Lines of Code Tokens Datapoints
@@
Simple 6 108 25.6 8 212 47.2 9 18 13.4 1 1 1.0 6 18 8.6 18 48 26.5 17
Moderate 6 344 78.6 8 684 155.0 9 87 30.6 1 5 2.2 7 21 12.6 22 60 35.2 39
Complex 8 540 93.1 12 1092 184.8 13 104 53.9 1 7 3.6 7 32 16.9 25 82 47.3 87
Very Complex 12 531 108.0 20 1070 214.4 13 188 68.5 1 11 4.5 6 46 21.3 18 117 53.6 57
All 6 540 88.8 8 1092 175.7 9 188 50.13 1 11 3.4 6 46 16.6 18 117 45.0 200